Regression equation for Case 3.0:

SUMMARY OUTPUT

Regression Statistics

Multiple R

0.957

R Square

0.915

Adjusted R Square

0.908

Standard Error

5.779

Observations

52

ANOVA

df

SS

MS

F

Significance F

Regression

4

16947.86487

4236.9662

126.8841

1.45976E-24

Residual

47

1569.442824

33.392401

Total

51

18517.30769

Coefficients

Standard Error

t Stat

P-value

Lower 95%

Upper 95%

Intercept

39.08190

15.31261

2.55227

0.014012

8.27693

69.88687

X-Price

-7.37039

0.98942

-7.44921

1.71E-09

-9.36084

-5.37994

Y-Price

-3.42813

0.21342

-16.06289

1.03E-20

-6.10796

-4.74831

Z-Price

4.05067

0.33949

11.93173

7.95E-16

3.36771

4.73363

Income

0.00288

0.00038

7.57448

1.11E-09

0.00212

0.00364

Questions and analysis:

5.    Suppose the marginal cost of model X is a constant $5 per unit. Find the profit maximizing price and quantity for the producer of model X, once again assuming the price of Y is $15, the price of Z is $24, and household income is $42,000.

Optimal Price: __________

Optimal Quantity: ___________

6.    Calculate cross price elasticity between the model X and the price of Y when own price is $10, the price of Y is $15, the price of Z is $24, and household income is $42,000. Suggest a strategic response: how exactly should the producer of model X respond when the producer of Y raises its price, say by $1?

Cross Price Elasticity = _______________

Strategic response:
Answer #1

The regression estimate would be Qx = 39.0819- 7.37039 PX - 3.42813PY + 4.05067P2 +0.002881 .

5. The marginal cost of X is MC=5. For the given prices and income, we have Qx = 39.0819-7.37039 PX-3.42813*15+4.05067* 24+0.00288*42000 or Qx = 205.83603- 7.37039 PX or Px = 27.927427178 -0.1356780310x as the (inverse) demand for X.

The total revenue would be TR=0xPx = 27.9274271780 x -0.1356780310 and the marginal revenue would be _ dặTR) MR= dox or d MR= (27.9274271780x -0.1356780310) dox or MR = 27.927427178-0.2713560620x .

The profit maximization would be where the MC is equal to the MR, ie where MC = MR or 5 = 27.927427178 -0.2713560620x or \widehat Q_X = 84.492039754 \approx 84.492 untis. The corresponding price would be P = 27.927427178 -0.13567 8031 * 84.492039754 or Px = 16.463713589 16.46 dollars.

Hence, the optimal price is $16.46 and the optimal quantity is 84.492 units of X.

6. For given price of X and Z, and income, we have \widehat Q_X = 39.0819 - 7.37039 *10 - 3.42813 P_Y + 4.05067 *24 + 0.00288*42000 or \widehat Q_X = 183.55408 - 3.42813 P_Y . The cross price elasticity between X and Y is e_c = \frac{\mathrm{d} \widehat Q_X / \widehat Q_X}{\mathrm{d} P_Y / P_Y} or e_c = \frac{P_Y}{\widehat Q_X} * \frac{\mathrm{d} \widehat Q_X}{\mathrm{d} P_Y} or e_c = \frac{P_Y}{\widehat Q_X} * \frac{\mathrm{d} }{\mathrm{d} P_Y}(183.55408 - 3.42813 P_Y) or HETSTHE-)* or e_c = (- 3.42813)\frac{P_Y}{\widehat Q_X} .

For price of Y be 15, the quantity would be \widehat Q_X = 183.55408 - 3.42813 *15 = 132.13213 . The cross price elasticity would hence be e_c = (- 3.42813)\frac{15}{132.13213} or e_c = - 0.389170673 \approx - 0.3892 . This means that for a marginal unit percent increase in price of Y, the quantity of X would reduce by 0.3892%.

The stratergic response would depend on the found cross price elasticity. For a rise in price of Y by $1 would mean that the price of Y increased by \frac{1}{15}*100 = 6.666666667 percent. The demand for X would reduce by 0.389170673*6.666666667 = 2.594471153 \approx 2.5945 percent. As demand reduces, the producer must increase the price of X.

The quantity would hence reduce to 132.1321(1 - \frac{2.5945}{100}) = 128.703932665 , and since Qx = 205.83603- 7.37039 PX , we have the required price as Px = 10.465131063 10.47 dollars.

Hence, cross price elasticity is -0.3892 and the strategic response would be to to increase the price by $0.47.